# ChatSpamDetector: Leveraging Large Language Models for Effective Phishing Email Detection

### v1

@@ Abstract
The proliferation of phishing sites and emails poses significant challenges to existing cybersecurity efforts. Despite advances in spam filters and email security protocols, problems with oversight and false positives persist. Users often struggle to understand why emails are flagged as spam, risking the possibility of missing important communications or mistakenly trusting phishing emails.
@@ 1 Introduction
The damage caused by phishing sites is growing every year, with phishing emails being actively used by attackers as a longstanding method of luring users. Modern email services have incorporated sophisticated spam filters and authentication protocols to mitigate these threats by identifying suspicious senders and analyzing email content for potential phishing indicators [17, 34]. However, despite these efforts, oversights and misclassifications of phishing emails persist. Many email services may classify emails as spam without providing users with an explicit rationale for the classification decision. As a result, recipients cannot determine if an email is truly phishing and may accidentally miss an important message or unintentionally click on a malicious link.
In this study, we propose a system, ChatSpamDetector, that uses Large Language Models (LLMs), such as ChatGPT, to effectively detect phishing emails. By generating prompts to LLMs based on email body and header information, our system leverages the powerful analytical capabilities of LLMs to detect phishing emails. This approach not only enables highly accurate identification of phishing emails, but also provides detailed explanations for the classification decision. By automatically generating reports detailing the classification results, our system helps users make informed decisions for subsequent actions, such as discarding suspicious emails or trusting legitimate ones. To evaluate the performance of our proposed system, we prepared a dataset and compared the detection accuracy with several LLMs and baseline systems. The results showed that our system using GPT-4 achieved an accuracy of 99.70%, outperforming other models and baseline systems. Using our unique prompts, this system increases the true positive rate compared to simple prompts. Through detailed analysis of LLM responses, we confirmed the ability of LLMs to extract key indicators from the headers and body of emails, prioritize them, and generate accurate responses, confirming their effectiveness in phishing detection. We have observed that the proposed system accurately identifies sender spoofing and brand impersonation, along with various social engineering (SE) techniques (e.g., account anomaly notifications, cryptocurrency airdrops, virus infection alerts) contained in phishing emails. The system can also reveal the various techniques attackers use to evade spam filters.
@@
  • We conducted evaluation experiments on datasets containing recent phishing emails and demonstrated that our proposed system using GPT-4 achieved 99.70% accuracy, outperforming other models and baseline systems.
@@ 2 Background
To enhance defenses against phishing emails, many studies have been proposed that use machine learning-based methods to analyze and categorize emails. These methods statistically extract phishing-specific features from headers and bodies for binary (ham/spam) classification [10, 35, 36, 45]. In addition, methods using natural language processing to examine the text of phishing emails and interpret their context and intent have been explored[12, 31]. These methods range from simple frequency-based feature vector generation, such as TF-IDF [5] and Bag-of-words [6], to advanced deep learning models such as BERT (Bidirectional Encoder Representations from Transformers), which use semantic analysis to detect SE techniques [30, 38, 40]. While many of these machine learning-based methods demonstrate high performance on existing spam email datasets, they primarily offer statistical detection results without providing specific rationales. Furthermore, models trained on limited datasets may not perform effectively against various types of new phishing emails in different languages.
@@ 3 ChatSpamDetector
In this study, we present ChatSpamDetector, a system that uses LLMs to detect phishing emails. Our system converts emails into appropriate prompts for LLMs to analyze the entire email content. This enables advanced contextual analysis to detect complex SE techniques that cannot be detected by previous methods, including advanced deep learning techniques. Even for phishing emails that bypass email authentication, the system can identify brand impersonation based on the difference between the brand name in the email and its legitimate domain name. In addition, unlike spam filters embedded in email services, our system not only outputs the detection results, but also provides a detailed report, including the rationale behind its decisions. Our system accepts email data in .eml format and outputs a JSON report containing the phishing detection results. Targeted phishing emails include those that contain links to phishing sites designed to trick recipients into visiting them. The process of our system—email parsing and decoding, simplification, prompt generation, and request handling—is detailed below.
@@ 3.1 Parsing and Decoding Email
An .eml file represents an email message in text format that includes a header containing metadata and formatting information, such as sender and time stamp information, and a body containing the message and attached files. Since headers and bodies can be encoded in different character sets and encoding schemes, our system decodes them and reconstructs the email for easier interpretation by LLMs. To avoid unnecessary overhead in LLM analysis, header information such as DKIM and DMARC authentication signatures or custom headers starting with X- are removed. For experimental fairness, headers related to any spam filtering decisions made by email servers are excluded from our dataset. Attachments are handled by extracting filenames and discarding binary data.
@@ 3.2 Simplifying Email
Prompt Template 1: Normal Prompt I want you to act as a spam detector to determine whether a given email is a phishing email or a legitimate email. Your analysis should be thorough and evidence-based. Phishing emails often impersonate legitimate brands and use social engineering techniques to deceive users. These techniques include, but are not limited to: fake rewards, fake warnings about account problems, and creating a sense of urgency or interest. Spoofing the sender address and embedding deceptive HTML links are also common tactics. Analyze the email by following these steps: 1. Identify any impersonation of well-known brands. 2. Examine the email header for spoofing signs, such as discrepancies in the sender name or email address. Evaluate the subject line for typical phishing characteristics (e.g., urgency, promise of reward). Note that the To address has been replaced with a dummy address. 3. Analyze the email body for social engineering tactics designed to induce clicks on hyperlinks. Inspect URLs to determine if they are misleading or lead to suspicious websites. 4. Provide a comprehensive evaluation of the email, highlighting specific elements that support your conclusion. Include a detailed explanation of any phishing or legitimacy indicators found in the email. 5. Summarize your findings and provide your final verdict on the legitimacy of the email, supported by the evidence you gathered. Email: ‘‘‘<Insert email text data here>’’’
@@ 3.4 Requesting
is_phishing boolean
A boolean value indicating whether the email is phishing (true) or legitimate (false).
phishing_score number
Phishing risk confidence score as an integer on a scale from 0 to 10.
brand_impersonated string
Brand name associated with the email, if applicable.
rationales string
Detailed rationales for the determination, up to 500 words.
brief_reason string
Brief reason for the determination.
@@ 5.1 Model Selection
Prompt Template 2: Simple Prompt Determine whether a given email is a phishing email or a legitimate email. Email: ‘‘‘<Insert email text data here>’’’
Prompt Template 3: Additional Prompt for LLama 2 (Note: Here is the same as Prompt Template 3.2, except for the email insertion part.) 6. Your output should be JSON-formatted text with the following keys: - is_phishing: a boolean value indicating whether the email is phishing (true) or legitimate (false) - phishing_score: phishing risk confidence score as an integer on a scale from 0 to 10 - brand_impersonated: brand name associated with the email, if applicable - rationales: detailed rationales for the determination, up to 500 words - brief_reason: brief reason for the determination Email: ‘‘‘<Insert email text data here>’’’
@@ 5.2 Results
Summary First, we measured recall, also known as the true positive rate ($=\frac{TP}{TP+FN}$), where TP represents true positives and FN represents false negatives. GPT-4 with the normal prompt was the highest at 99.70%, outperforming GPT-3.5 (97.03%), Llama2 (94.06%) and Gemini Pro (98.12%). Comparisons of results between simple and normal prompts show an improvement in recall across all models, indicating that simple prompts tend to be significantly biased towards negative (legitimate) classifications. In fact, the false positives were significantly lower with simple prompts across all models, with 4, 6, 9, and 6 for GPT-4, GPT-3.5, Llama 2, and Gemini Pro, respectively. The use of normal prompts improved the models’ ability to accurately identify phishing emails while reducing false positives. This improvement was particularly pronounced for the less performing models, with GPT-3.5 and Llama2 having recalls increased by 28.02% and 15.84%, respectively. However, the normal prompt for Llama 2 caused an increase in false positives (FPs) in contrast to the simple prompts, leading to a 26.41% reduction in precision ($=\frac{TP}{TP+FP}$). When evaluating precision and accuracy ($=\frac{TP+TN}{TP+FP+TN+FN}$, where TN represents true negatives), both metrics for GPT-4 using normal prompts were exceptionally high at 99.70%, demonstrating its superiority over the other models.
@@
Phishing Score Figures 0(a) and 0(b) show the distribution of the phishing_score when using GPT-4 and GPT-3.5 respectively, categorized into TP, TN, FP and FN with stacked bars. Overall, emails classified as phishing had higher scores, while those classified as legitimate had lower scores. For GPT-4, the scores for emails classified as legitimate (negative) ranged from 1 to 4, with TN scoring most frequently 0 (604 cases), indicating a high level of confidence in correctly identifying legitimate emails. On the other hand, scores for emails identified as phishing (positive) were distributed between 6 and 10, with peaks at 8 and 9. GPT-3.5 had high scores such as 8 and 9 for FP and low scores such as 0 for FN, while GPT-4’s FP and FN were more neutrally distributed around 6, 7, 8 for FP and 1, 2 for FN, indicating that GPT-4 is less likely to make mistakes when providing higher confidence responses.
API Request Latency We describe the response time when experimenting with the proposed system. The average time from sending the request to receiving the response was 12.53 seconds for GPT-4 and 2.50 seconds for GPT-3.5. Note that these results are subject to change based on various factors, such as the region of the Azure OpenAI Service and server congestion.
Cost Analysis We did a cost analysis for the experiments using the ChatGPT API (GPT-4 and GPT-3.5). The cost per input was \$0.03 for GPT-4 and \$0.002 for GPT-3.5, while the cost per output was \$0.06 for GPT-4 and \$0.002 for GPT-3.5. Analyzing the 2,010 emails in our dataset, the cost of using GPT-4 with a normal prompt was \$265.80 and for GPT-3.5 was \$13.68, while using a simple prompt resulted in costs of \$249.93 for GPT-4 and \$12.62 for GPT-3.5.
@@ 5.4 Examples of Detection Results
Response 1: Example A { "is_phishing": true, "phishing_score": 8, "brand_impersonated": "Microsoft", "rationales": "This email appears to be a phishing attempt. The sender is ostensibly from ’crous-montpellier.fr’, but the email’s body text mentions the ’Microsoft account team’, which is an impersonation of the Microsoft brand. This discrepancy is a sign of phishing. The email subject line, ’Your new user agreement’, is designed to create a sense of urgency and lure the recipient into clicking on the link provided. The hyperlink uses a tracking ID which suggests that it might be used for collecting sensitive information. Examining the authentication results from the header, the emails passed SPF (Sender Policy Framework) and DKIM (DomainKeys Identified Mail) checks. However, this could still be sophisticated phishing, as attackers nowadays often compromise legitimate servers to launch phishing campaigns.", "brief_reason": "Impersonation of a well-known brand (Microsoft), inconsistency in sender’s information, and presence of a suspicious hyperlink." }
Response 2: Example B { "is_phishing": true, "phishing_score": 9, "brand_impersonated": "Binance", "rationales": "Several indicators lean towards this being a phishing email. To begin with, the sender’s email address has a discrepancy: the email claims to be from ’Binance’, but the sender’s domain is not associated with Binance (ilonasavola.com). Moreover, the SPF check has failed as indicated in the ’Authentication-Results’ field. The content of mail also raises suspicion: it impersonates Binance, creates a sense of urgency by stating that withdrawals are disabled and the account will be permanently disabled without immediate action and uses a disguised link for ’update informations’, which leads to a different domain (zzdzw.com). These elements are common tactics employed in phishing emails, designed to deceive users into providing personal information.", "brief_reason": "Spoofed sender address and domain, impersonation of Binance, sense of urgency, disguised potentially harmful link, and SPF check failure." }
@@ 5.5 LLM Capabilities for Phishing Email Detection
Extracting and Comparing Information Contained in Headers LLMs perform advanced analysis of email headers, extracting critical information primarily from the sender address, authentication results (pass/fail), subject line, and communication paths. The sender’s address is the most important factor in analyzing the authenticity of an email. Although the From header is easily spoofed, LLMs can determine the legitimacy of an email by combining the address with the authentication results. In addition, LLMs check for discrepancies between the sender’s domain name and the originating server in the Received header. For example, GPT-4 flagged a phishing email purportedly from Trust Wallet <noreply@support-trustwallet[.]com> as a spoofing attempt based on two critical observations: the domain name did not appear in the Received line, and the Authentication-Results field indicated a softfail for both SPF and DMARC checks. LLMs also identify SE techniques from the subject line, a key factor in enticing users to open an email, including psychological tactics to create curiosity or a sense of urgency.

### v2

@@ Abstract
The proliferation of phishing sites and emails poses significant challenges to existing cybersecurity efforts. Despite advances in malicious email filters and email security protocols, problems with oversight and false positives persist. Users often struggle to understand why emails are flagged as potentially fraudulent, risking the possibility of missing important communications or mistakenly trusting deceptive phishing emails.
@@
If you cite this paper, please use the following reference:
Takashi Koide, Naoki Fukushi, Hiroki Nakano, and Daiki Chiba.: ChatSpamDetector: Leveraging Large Language Models for Effective Phishing Email Detection. In 20th EAI International Conference on Security and Privacy in Communication Networks (SecureComm 2024), October 28–30, 2024, Dubai, United Arab Emirates.
@@ 1 Introduction
The damage caused by phishing sites is growing every year, with phishing emails being actively used by attackers as a longstanding method of luring users. Modern email services have incorporated sophisticated malicious email filters and authentication protocols to mitigate these threats by identifying suspicious senders and analyzing email content for potential phishing indicators [18, 36]. However, despite these efforts, oversights and misclassifications of phishing emails persist. Many email services may classify emails as potentially fraudulent without providing users with an explicit rationale for the classification decision. As a result, recipients cannot determine if an email is truly phishing and may accidentally miss an important message or unintentionally click on a malicious link.
In this study, we propose a system, ChatSpamDetector, that uses Large Language Models (LLMs), such as ChatGPT, to effectively detect phishing emails. By generating prompts to LLMs based on email body and header information, our system leverages the powerful analytical capabilities of LLMs to detect phishing emails. This approach not only enables highly accurate identification of phishing emails, but also provides detailed explanations for the classification decision. While previous research [22] has explored the use of LLMs to automatically generate phishing emails and analyze their intent, the ability of LLMs to distinguish between phishing and legitimate emails has not been clearly demonstrated. Our system is the first study to evaluate the ability of LLMs to detect phishing emails.
By automatically generating reports detailing the classification results, our system helps users make informed decisions for subsequent actions, such as discarding suspicious emails or trusting legitimate ones. To evaluate the performance of our proposed system, we prepared a dataset containing both phishing emails and legitimate emails, and compared the detection accuracy with several LLMs and baseline systems. The results showed that our system using GPT-4 achieved an accuracy of 99.70%, outperforming other models and baseline systems. Using our unique prompts, this system increases the true positive rate compared to simple prompts.
Through detailed analysis of LLM responses, we confirmed the ability of LLMs to extract key indicators from the headers and body of emails, prioritize them, and generate accurate responses, confirming their effectiveness in phishing detection. We have observed that the proposed system accurately identifies sender spoofing and brand impersonation, along with various social engineering (SE) techniques (e.g., account anomaly notifications, cryptocurrency airdrops, virus infection alerts) contained in phishing emails. The system can also reveal the various techniques attackers use to evade malicious email filters.
Our system represents a new step in the fight against phishing emails, as it can be used to replace or complement the existing phishing detection features of email services, helping users to determine the suspiciousness of emails on their own. By providing users with the tools to make informed decisions about the legitimacy of emails, our proposed system enables individuals to protect themselves against the evolving phishing attack landscape.
@@
  • We collected recent phishing emails and legitimate emails to create datasets, and conducted evaluation experiments on them. The results demonstrated that our proposed system using GPT-4 achieved 99.70% accuracy, outperforming other models and baseline systems.
@@ 2 Background
To enhance defenses against phishing emails, many studies have been proposed that use machine learning-based methods to analyze and categorize emails. These methods statistically extract phishing-specific features from headers and bodies for binary (ham/spam) classification [10, 37, 38, 48]. In addition, methods using natural language processing to examine the text of phishing emails and interpret their context and intent have been explored[13, 33]. These methods range from simple frequency-based feature vector generation, such as TF-IDF [5] and Bag-of-words [6], to advanced deep learning models such as BERT (Bidirectional Encoder Representations from Transformers), which use semantic analysis to detect SE techniques [32, 40, 42]. While many of these machine learning-based methods demonstrate high performance on existing phishing email datasets, they primarily offer statistical detection results without providing specific rationales. Furthermore, models trained on limited datasets may not perform effectively against various types of new phishing emails in different languages.
@@ 3 ChatSpamDetector
Figure 1: System Architecture of ChatSpamDetector
In this study, we present ChatSpamDetector, a system that uses LLMs to detect phishing emails. Our system converts emails into appropriate prompts for LLMs to analyze the entire email content. This enables advanced contextual analysis to detect complex SE techniques that cannot be detected by previous methods, including advanced deep learning techniques. Even for phishing emails that bypass email authentication, the system can identify brand impersonation based on the difference between the brand name in the email and its legitimate domain name. In addition, unlike spam filters embedded in email services, our system not only outputs the detection results, but also provides a detailed report, including the rationale behind its decisions. Our system accepts email data in .eml format and outputs a JSON report containing the phishing detection results. Targeted phishing emails include those that contain links to phishing sites designed to trick recipients into visiting them. The system architecture of ChatSpamDetector is shown in Figure 1. The process of our system—email parsing and decoding, simplification, prompt generation, and request handling—is detailed below.
@@ 3.1 Parsing and Decoding Email
An .eml file represents an email message in text format that includes a header containing metadata and formatting information, such as sender and time stamp information, and a body containing the message and attached files. Since headers and bodies can be encoded in different character sets and encoding schemes, our system decodes them and reconstructs the email for easier interpretation by LLMs. For example, the email subject Subject: Security Alert! when encoded using the UTF-8 character set and Base64 encoding, appears as Subject: =?UTF-8?B?U3ViamVjdDogU2VjdXJpdHkgQWxlcnQh?=. Such strings, while generally interpretable by the advanced parsing capabilities of LLMs for short texts, increasingly pose a risk of failure as they lengthen.
To avoid unnecessary overhead in LLM analysis, header information such as DKIM and DMARC authentication signatures or custom headers starting with X- are removed. For experimental fairness, headers related to any spam filtering decisions made by email servers are excluded from our dataset. Attachments are handled by extracting filenames and discarding binary data.
@@ 3.2 Simplifying Email
Prompt Template 1: Normal Prompt I want you to act as a spam detector to determine whether a given email is a phishing email or a legitimate email. Your analysis should be thorough and evidence-based. Phishing emails often impersonate legitimate brands and use social engineering techniques to deceive users. These techniques include, but are not limited to: fake rewards, fake warnings about account problems, and creating a sense of urgency or interest. Spoofing the sender address and embedding deceptive HTML links are also common tactics.
Analyze the email by following these steps:
1. Identify any impersonation of well-known brands.
2. Examine the email header for spoofing signs, such as discrepancies in the sender name or email address.
Evaluate the subject line for typical phishing characteristics (e.g., urgency, promise of reward). Note that the To address has been replaced with a dummy address.
3. Analyze the email body for social engineering tactics designed to induce clicks on hyperlinks. Inspect URLs to determine if they are misleading or lead to suspicious websites.
4. Provide a comprehensive evaluation of the email, highlighting specific elements that support your conclusion. Include a detailed explanation of any phishing or legitimacy indicators found in the email.
5. Summarize your findings and provide your final verdict on the legitimacy of the email, supported by the evidence you gathered.
Email:
‘‘‘<Insert email text data here>’’’
@@ 3.4 Requesting
is_phishing boolean A boolean value indicating whether the email is phishing (true) or legitimate (false).
phishing_score number Phishing risk confidence score as an integer on a scale from 0 to 10.
brand_impersonated string Brand name associated with the email, if applicable.
rationales string Detailed rationales for the determination, up to 500 words.
brief_reason string Brief reason for the determination.
@@ 5.1 Model Selection
Prompt Template 2: Simple Prompt Determine whether a given email is a phishing email or a legitimate email.
Email:
‘‘‘<Insert email text data here>’’’
Prompt Template 3: Additional Prompt for LLama 2 Note: Here is the same as Prompt Template 3.2, except for the email insertion part. 6. Your output should be JSON-formatted text with the following keys:
- is_phishing: a boolean value indicating whether the email is phishing (true) or legitimate (false)
- phishing_score: phishing risk confidence score as an integer on a scale from 0 to 10
- brand_impersonated: brand name associated with the email, if applicable
- rationales: detailed rationales for the determination, up to 500 words
- brief_reason: brief reason for the determination
Email:
‘‘‘<Insert email text data here>’’’
@@ 5.2 Results
5.2.1 Summary
First, we measured recall, also known as the true positive rate ($=\frac{TP}{TP+FN}$), where TP represents true positives and FN represents false negatives. GPT-4 with the normal prompt was the highest at 99.70%, outperforming GPT-3.5 (97.03%), Llama2 (94.06%) and Gemini Pro (98.12%). Comparisons of results between simple and normal prompts show an improvement in recall across all models, indicating that simple prompts tend to be significantly biased towards negative (legitimate) classifications. In fact, the false positives were significantly lower with simple prompts across all models, with 4, 6, 9, and 6 for GPT-4, GPT-3.5, Llama 2, and Gemini Pro, respectively. The use of normal prompts improved the models’ ability to accurately identify phishing emails while reducing false positives. This improvement was particularly pronounced for the less performing models, with GPT-3.5 and Llama2 having recalls increased by 28.02% and 15.84%, respectively. However, the normal prompt for Llama 2 caused an increase in false positives (FPs) in contrast to the simple prompts, leading to a 26.41% reduction in precision ($=\frac{TP}{TP+FP}$). When evaluating precision and accuracy ($=\frac{TP+TN}{TP+FP+TN+FN}$, where TN represents true negatives), both metrics for GPT-4 using normal prompts were exceptionally high at 99.70%, demonstrating its superiority over the other models.
@@ 5.2.1 Summary
5.2.2 Phishing Score
Figures 2(a) and 2(b) show the distribution of the phishing_score when using GPT-4 and GPT-3.5 respectively, categorized into TP, TN, FP and FN with stacked bars. Overall, emails classified as phishing had higher scores, while those classified as legitimate had lower scores. For GPT-4, the scores for emails classified as legitimate (negative) ranged from 1 to 4, with TN scoring most frequently 0 (604 cases), indicating a high level of confidence in correctly identifying legitimate emails. On the other hand, scores for emails identified as phishing (positive) were distributed between 6 and 10, with peaks at 8 and 9. GPT-3.5 had high scores such as 8 and 9 for FP and low scores such as 0 for FN, while GPT-4’s FP and FN were more neutrally distributed around 6, 7, 8 for FP and 1, 2 for FN, indicating that GPT-4 is less likely to make mistakes when providing higher confidence responses.
5.2.3 API Request Latency
We describe the response time when experimenting with the proposed system. The average time from sending the request to receiving the response was 12.53 seconds for GPT-4 and 2.50 seconds for GPT-3.5. Note that these results are subject to change based on various factors, such as the region of the Azure OpenAI Service and server congestion.
5.2.4 Cost Analysis
We did a cost analysis for the experiments using the ChatGPT API (GPT-4 and GPT-3.5). The cost per input was \$0.03 for GPT-4 and \$0.002 for GPT-3.5, while the cost per output was \$0.06 for GPT-4 and \$0.002 for GPT-3.5. Analyzing the 2,010 emails in our dataset, the cost of using GPT-4 with a normal prompt was \$265.80 and for GPT-3.5 was \$13.68, while using a simple prompt resulted in costs of \$249.93 for GPT-4 and \$12.62 for GPT-3.5.
@@ 5.4 Examples of Detection Results
Response 1: Example A {
"is_phishing": true,
"phishing_score": 8,
"brand_impersonated": "Microsoft",
"rationales": "This email appears to be a phishing attempt. The sender is ostensibly from ’crous-montpellier.fr’, but the email’s body text mentions the ’Microsoft account team’, which is an impersonation of the Microsoft brand. This discrepancy is a sign of phishing. The email subject line, ’Your new user agreement’, is designed to create a sense of urgency and lure the recipient into clicking on the link provided. The hyperlink uses a tracking ID which suggests that it might be used for collecting sensitive information. Examining the authentication results from the header, the emails passed SPF (Sender Policy Framework) and DKIM (DomainKeys Identified Mail) checks. However, this could still be sophisticated phishing, as attackers nowadays often compromise legitimate servers to launch phishing campaigns.",
"brief_reason": "Impersonation of a well-known brand (Microsoft), inconsistency in sender’s information, and presence of a suspicious hyperlink."
Response 2: Example B {
"is_phishing": true,
"phishing_score": 9,
"brand_impersonated": "Binance",
"rationales": "Several indicators lean towards this being a phishing email. To begin with, the sender’s email address has a discrepancy: the email claims to be from ’Binance’, but the sender’s domain is not associated with Binance (ilonasavola.com). Moreover, the SPF check has failed as indicated in the ’Authentication-Results’ field. The content of mail also raises suspicion: it impersonates Binance, creates a sense of urgency by stating that withdrawals are disabled and the account will be permanently disabled without immediate action and uses a disguised link for ’update informations’, which leads to a different domain (zzdzw.com). These elements are common tactics employed in phishing emails, designed to deceive users into providing personal information.",
"brief_reason": "Spoofed sender address and domain, impersonation of Binance, sense of urgency, disguised potentially harmful link, and SPF check failure."
@@ 5.5 LLM Capabilities for Phishing Email Detection
5.5.1 Extracting and Comparing Information Contained in Headers
LLMs perform advanced analysis of email headers, extracting critical information primarily from the sender address, authentication results (pass/fail), subject line, and communication paths. The sender’s address is the most important factor in analyzing the authenticity of an email. Although the From header is easily spoofed, LLMs can determine the legitimacy of an email by combining the address with the authentication results. In addition, LLMs check for discrepancies between the sender’s domain name and the originating server in the Received header. For example, GPT-4 flagged a phishing email purportedly from Trust Wallet <noreply@support-trustwallet[.]com> as a spoofing attempt based on two critical observations: the domain name did not appear in the Received line, and the Authentication-Results field indicated a softfail for both SPF and DMARC checks. LLMs also identify SE techniques from the subject line, a key factor in enticing users to open an email, including psychological tactics to create curiosity or a sense of urgency.
5.5.2 Identifying SE Techniques in the Body
The body of an email often contains various SE techniques aimed at encouraging users to click on links. Phishing emails typically exploit the trust and recognition of well-known brands to lower users’ caution and encourage them to follow the instructions. Attackers try to deceive recipients by constructing convincing scenarios, such as “payment confirmation” posing as online banking or “security issues with your account” posing as a social media platform. In these cases, LLMs can accurately identify the brand impersonation, not merely through Named Entity Recognition (NER) but by analyzing the context to determine which brand is being impersonated.